OpenStudy (anonymous):

h(x)=-3x^2-6 find h(x+5)

OpenStudy (anonymous):

Simply substitute the x in h(x) with x+5 instead: h(x+5)=-3(x+5)^2-6

OpenStudy (anonymous):

so solve -3(x+5)^2-6 ??

OpenStudy (anonymous):

What do you mean by solve? The question doesn't ask you to solve for anything.

OpenStudy (anonymous):

so gven h(x+5)=-3(x+5)^2-6 what would be the next thing i do square the x+5 ?

OpenStudy (anonymous):

If you want, yes. That would just be further simplifying h(x+5) after you've already found it.

OpenStudy (anonymous):

so what would i do with the h(x+5) on the left hand side ?

OpenStudy (anonymous):

this would be -3(x^2+10x+25)-6 which is -3x^2-30x-81, end.

OpenStudy (anonymous):

Nothing. The whole point is to get h(x+5) since your questions asks us to "find h(x+5)."

OpenStudy (anonymous):

nothing, just leave it, this is what h(x+5) is equal to

OpenStudy (anonymous):

okay thanks